OpenStudy (anonymous):

Find the inverse of the function. f(x) = the cube root of quantity x divided by seven. - 9 answers: f-1(x) = 21(x + 9) f-1(x) = [7(x + 9)]3 f-1(x) = 7(x3 + 9) f-1(x) = 7(x + 9)3

OpenStudy (agent0smith):

Can you write it as an equation, not words?

OpenStudy (anonymous):

\[\sqrt[3]{x/7}-9\]

OpenStudy (agent0smith):

\[\huge y = \sqrt[3]{ \frac{ x }{ 7 } } -9\]solve this for y. Start by adding 9 to both sides. Then you'll have to cube both sides, to get rid of the cube root.

OpenStudy (anonymous):

okay I did that, I don't know how it turns out when you cube it, that parts messing me up

OpenStudy (agent0smith):

\[\huge y +9 = \sqrt[3]{ \frac{ x }{ 7 } } \]cubing it does this:\[\huge (y+9)^3 = \frac{ x }{ 7 } \]

OpenStudy (anonymous):

oh okay is it, \[f^1(x)=21(x+9)\]?

OpenStudy (agent0smith):

No. All you have to do is multiply both sides by 7, then swap x and y.

OpenStudy (agent0smith):

Leave the cubed part alone.